undeadknight26 (undeadknight26):

Help to convert to standard equation? o f(x) = x^2 + 6x – 16 o g(x) = x^2 +6x + 1

undeadknight26 (undeadknight26):

@iGreen @ganeshie8

undeadknight26 (undeadknight26):

So far i have: 1. To open these doors, you must match the number and type of solutions for the following two functions in standard form. o f(x) = x2 + 6x – 16 x^2 + 8x – 2x – 16 (x – 8)-2(x-8) (x+8)(x-2) So the solutions would be -8 and 2. o g(x) = x2 +6x + 1 x^2 + x8 – 2x + 1 x(x + 8) -2x + 1 So the solutions are -8 and 2. Is it correct?

undeadknight26 (undeadknight26):

@Yahoo! @YanaSidlinskiy @midhun.madhu1987 ?

OpenStudy (yanasidlinskiy):

That's what I got.

OpenStudy (emogirlatemoooocow):

so it would be (6)^2-4(1)(16) Simplify that

OpenStudy (midhun.madhu1987):

f(x) is fine.. but g(x) cant be simplified this way... you need use Quadratic Formula .. i guess...

undeadknight26 (undeadknight26):

Can you help me with it lel? i kinda suck at these things ._.

OpenStudy (yanasidlinskiy):

\(y = ax^2 + bx + c\) We all suck at it!!

OpenStudy (emogirlatemoooocow):

@undeadknight26 it would be -8 right?.

OpenStudy (emogirlatemoooocow):

*-28

OpenStudy (emogirlatemoooocow):

i take tht back it 64

undeadknight26 (undeadknight26):

o g(x) = x2 +6x + 1 x^2 + 8x - 2x + 1 x(x + 8)(x + 2) + 1

OpenStudy (emogirlatemoooocow):

so then it would be 36+64 which would be 100

undeadknight26 (undeadknight26):

I just need to put it into standard ._. I dunno what im doing ._.

OpenStudy (midhun.madhu1987):

Is the question completely written above?

undeadknight26 (undeadknight26):

o g(x) = x^2 +6x + 1 This is the G(x) equation.

OpenStudy (midhun.madhu1987):

I mean... is the question this much?? Help to convert to standard equation? or anything else is there?

undeadknight26 (undeadknight26):

Well there are other parts but for now i just need help with this one.

OpenStudy (zpupster):

http://www.mathwarehouse.com/geometry/parabola/standard-to-vertex-form.php standard form of parabola is ax^2+bx+c vertex stand form is y=a(x-h)^2 + k

OpenStudy (zpupster):

your equations are in standard form

undeadknight26 (undeadknight26):

They are?

undeadknight26 (undeadknight26):

g(x) = x2 +6x + 1 x^2 + 8x - 2x + 1 x(x + 8)(x + 2) + 1 (x+8)(x+2) Are u sure its correct?

OpenStudy (zpupster):

you can put them in vertex form complete the square and add the constants together follow the link i gave

undeadknight26 (undeadknight26):

I just need to put them in standard form for now ._. I ask again ._. Is it correct?!